# Enabling AI Scientists to Recognize Innovation: A Domain-Agnostic Algorithm for Assessing Novelty

### v2

@@ Abstract
In the pursuit of Artificial General Intelligence (AGI), automating the generation and evaluation of novel research ideas is a key challenge in AI-driven scientific discovery. This paper presents Relative Neighbor Density (RND), a domain-agnostic algorithm for novelty assessment in research ideas that overcomes the limitations of existing approaches by comparing an idea’s local density with its adjacent neighbors’ densities. We first developed a scalable methodology to create test set without expert labeling, addressing a fundamental challenge in novelty assessment. Using these test sets, we demonstrate that our RND algorithm achieves state-of-the-art (SOTA) performance in computer science (AUROC=0.820) and biomedical research (AUROC=0.765) domains. Most significantly, while SOTA models like Sonnet-3.7 and existing metrics show domain-specific performance degradation, RND maintains consistent accuracies across domains by its domain-invariant property, outperforming all benchmarks by a substantial margin (0.795 v.s. 0.597) on cross-domain evaluation. These results validate RND as a generalizable solution for automated novelty assessment in scientific research.
@@ 1 Introduction
In the pursuit of Artificial General Intelligence (AGI), automating scientific research and knowledge discovery presents both a formidable challenge and an exciting opportunity, as it will be groundbreaking to expand the boundaries of human knowledge by leveraging scalable computing resources. Therefore, as the capabilities of large language models (LLMs) continue to improve, researchers have started to explore their use in automating various aspects of the research process, including the generation of novel ideas, as exemplified by the AI scientist concept (Lu et al. (2024)).
A key task for any AI-based scientist is the generation of novel research ideas, a task traditionally performed by human scientists during their brainstorming phase. While LLMs have shown promise in generating a large pool of ideas quickly and cost-effectively, akin to the initial stages of human research, the real challenge lies in evaluating these ideas for their novelty. Traditionally, novelty in scientific research has been assessed through peer review and expert evaluations, where domain specialists judge the originality of an idea based on their experience and familiarity with existing literature. However, such assessments are inherently subjective, time-consuming, and inconsistent across reviewers. Moreover, as the volume of scientific output grows exponentially, manual novelty assessment struggles to keep pace. Automated methods are therefore crucial for filtering out redundant ideas and promoting genuinely innovative directions.
Existing approaches primarily fall into two categories: (1) leveraging large language models (LLMs) as judges and (2) using absolute local density-based novelty metrics.
The most straightforward approach is to use LLMs as judges to evaluate the novelty of ideas. Si. et al. adopted a Swiss system tournament design to evaluate ideas by using LLM as judge (Si et al. (2024)), which was further applied in Nova: An Iterative Planning and Search Approach to enhance Novelty and Diversity of LLM-Generated Ideas (Hu et al. (2024)). To improve LLM’s accuracy of judgment, Lu include NeurIPS review guideline and Semantic Scholar API as tools (Lu et al. (2024); Su et al. (2024)): the NeurIPS review guideline was served as both chain-of-thoughts prompts and few-shot examples, while search API enabled LLM to search top 10 relevant papers to determine the novelty of given idea.
An alternative approach relies on the absolute local density in semantic embedding space to measure novelty. Su. et al. introduced the concept of Historical Dissimilarity and Contemporary Dissimilarity, calculated as the average Euclidean distance (local density) between a generated abstract’s embedding and the embeddings of the five most similar abstracts from a historical database and a contemporary database, respectively. In combination with the citation-based Contemporary Impact metric, they developed the Overall Novelty (ON) metric(Su et al. (2024)). Their study validated that ON correlates well with human-labeled novelty, demonstrating its effectiveness as a novelty assessment measure.
Though the aforementioned approaches provide potential solutions on idea novelty evaluation, there are major challenges when considering practical issues.
First, the reliability of using-LLM-as-judge remains questionable, even with external knowledge or tools. Studies have demonstrated that auto-regressive LLMs like GPT-4o (Hurst et al. (2024)) produce outputs sensitive to input perturbations (Zhuo et al. (2023); Singh et al. (2024)). In novelty assessment specifically, this means identical research ideas phrased differently might receive contradictory novelty ratings. While recent reasoning models, such as DeepSeek-r1 (Guo et al. (2025)) and Sonnet-3.7 (Anthropic (2025)), show improved reasoning capabilities, their reliability for scientific novelty judgment remains unvalidated.
Second, the absolute local density-based metric from (Su et al. (2024)) shows significant limitations across diverse research contexts. By relying on just 5 most similar abstracts from history and contemporary databases as reference points, the metric’s validity becomes highly dependent on arbitrary choices: the size of paper collections, the temporal boundaries defining ’history’ versus ’contemporary,’ and the selection criteria for inclusion. Different research domains also exhibit varying citation patterns, publication velocities, and semantic densities, which undermined the metric’s generalizability across research domains.
Last but not least, the validation methodology used to assess novelty evaluators themselves was significantly lacking. In most cases, the validation of novelty metric depends on small test sets manually labeled by human experts within a specific domain (Hu et al. (2024); Lu et al. (2024); Su et al. (2024); Si et al. (2024)). Such validations are difficult to scale across different research areas, as they are often highly specialized and tailored to particular fields of study. What’s worse, manually produced novelty labels rapidly become outdated: as scientific research advances continuously, ideas labeled "novel" today quickly become established knowledge, rendering static human-labeled validation sets increasingly inaccurate over time.
@@
To address these challenges, in this paper, we establish comprehensive semantic embedding databases for novelty assessment. These databases incorporate over 30 million publications from two distinct domains: Pubmed, the leading biomedical literature search engine with nearly 36 million articles (Jin et al. (2024)), and Arxiv, which contains more than 2.3 million scholarly articles across eight subject areas (Cornell Tech (2023)).
@@
Based on these resources, we propose the Relative Neighbor Density (RND) algorithm, which measures novelty by analyzing the distribution patterns of semantic neighbors rather than simple absolute local density. This approach proves more reliable than LLM-based judgments and more generalizable than existing absolute local density-based metrics across different research domains. We also develop an automated validation methodology that leverages temporal publication patterns to evaluate novelty without requiring expert manual labeling. Our extensive evaluations using test sets from computer science, biomedical science, and cross-domain contexts demonstrate that our proposed algorithm maintains accuracy within specific domains while scaling effectively across diverse research areas.
@@
The remainder of this paper is organized as follows: Section 2 reviews related work in novelty assessment. Section 3 presents our methodology, including database construction, algorithmic details, and test set development. Section 4 describes our experimental results and validation across different domains. Finally, Section 5 discusses our findings, limitations, and future directions.
@@ 2 Related Works
Assessing the novelty of research ideas is a fundamental challenge in automating scientific discovery. Various approaches have been explored in recent years, leveraging Large Language Models (LLMs) and semantic similarity measures to evaluate idea originality. This section reviews existing methods, highlighting their strengths and limitations.
@@ 2.2 Absolute Local Density-based Metrics
An alternative approach focuses on semantic local density to evaluate novelty. Su et al. (Su et al. (2024)) used the Historical Dissimilarity (HD), which is the average Euclidean distance between the generated abstract embedding and embeddings of the 5 most similar abstracts in historical literature base. We denote it as "Absolute Local Density" because the average distance is a metric of local density, and they use the value of density directly. In addition to HD, they also used Contemporary Dissimilarity (CD), which calculate in same algorithm in contemporary literature base, which is also an absolute local density-based metric.
Based on HD and CD, they developed Overall Novelty (ON) metric as below,
@@ 2.3 Validating Novelty Metrics
Furthermore, the scalability challenge persists across all existing frameworks. The creation of test sets currently requires expert labeling, making large-scale evaluation prohibitively resource-intensive. This limitation underscores the need for automated approaches to test set creation that maintain fidelity while eliminating the requirement for extensive manual labeling.
@@ 3.2 Semantic Embedding & Literature Database
For each paper, two semantic embedding vectors were generated—one from its title and another from its abstract—using the M3-Embedding model Chen et al. (2024). The embedding vector dimension, denoted as $dims$, is 1024. All texts and embedding vectors were stored in Elasticsearch Version 8 for efficient retrieval.
@@ 3.3 Algorithm
The selection of values of $P$ and $Q$ is a trade-off between reliability of estimation and other cost. A lower values of $P$ will cause biased estimation of novelty score; while higher value of $P$ will increase computing cost in $O(P\cdot Q)$ complexity. Similarly, lower $Q$ will also cause unreliable estimation of local density. Meanwhile, a higher $Q$ value will not only be computational costly, but also sacrifice sensitivity as ND converges to its expectation as $Q$ increase, losing information local density. Refer to Appendix A for detailed analysis on the effects of $P$ and $Q$.
Based on analysis and empirical experiments, we set $P=100$ and $Q=50$.
The pseudoscope implement calculation of ND is presented in Algorithm 1; and the pseudoscope of complete algorithm is provided in Algorithm 2.
Algorithm 1 Find Neighbors and Calculate Neighbor Density
1:function Neighbor(Input, P, Q)
3:     C $\leftarrow$ []
5:     $neighbors\_for\_count\leftarrow neighbors[:Q]$triangleright Only use the Q nearest neighbors to calculate density
6:     $neighbors\_for\_distribution\leftarrow neighbors[:P]$triangleright The P nearest neighbors are used to calculate distribution
7:     for each paper in $neighbors\_for\_count$ do
8:         $\mathbf{v}_{paper}\leftarrow$ Get_Embedding(paper)
10:         C.Append(distance)
11:     end for
12:     $ND_{Input}\leftarrow$ mean(C)
13:     return $ND_{Input}$, $neighbors\_for\_distribution$
14:end function
Algorithm 2 Calculate Novelty Score of Given Idea
1:Input: Idea
2:Output: A score in the range of 0 to 100
3:D $\leftarrow$ []
4:$ND_{Idea}$, neighbors $\leftarrow$ Neighbor(Idea, P, Q)
5:for paper in neighbors do
6:     $ND_{paper}$, _ $\leftarrow$ Neighbor(paper, P, Q)
7:     D.Append($ND_{paper}$)
8:end for
9:score |$\leftarrow\frac{|\{ND\in D\mid ND\leq ND_{Idea}\}|}{|D|}\times 100$
10:Return score
@@ 3.4 Validation without Human Labeling
In this way, we can make positive and negative samples have a more obvious difference in novelty in a relatively objective and recognized way.
@@ 4.1.2 Baseline
To evaluate our algorithm, we selected all existing novelty assessment algorithms as baselines, categorized into two groups: LLM-based and non-LLM-based. Non-LLM-based algorithms, including Relative Neighbor Density(Ours), Historical Dissimilarity(HD), and Overall Novelty(ON), rely solely on literature search and mathematical calculations. Since the output of the literature search for the same query remains consistent, we conducted a single test to assess the algorithm’s performance. In contrast, for LLM-based algorithms, due to the inherent variability of LLM outputs, we ran three tests for each algorithm, calculated the average result, and included the standard deviation in the table. The full experimental results of the LLM-based method are provided in Table 9.
For all methods, we use the abstracts of the papers in the test set as "ideas" for testing.
Historical Dissimilarity: Identify the five most relevant papers based on their embeddings and compute the Euclidean distance between the embedding of the idea and the embeddings of the abstracts of these five papers. The final novelty score is obtained by averaging these distance values.
Overall Novelty: The historical database contains papers from 2011 to 2021, and the contemporary database contains papers from 2021 to 2025. The score calculation method refers to equation 1.
LLM + literature search: Provide LLM with the titles and abstracts of the 10 most relevant papers to the given idea. The model then assesses whether the core concepts of these papers significantly overlap with the idea(Table LABEL:tab:LLM_with_paper_prompt). If substantial overlap is detected, the idea is deemed non-novel and assigned a score of 0. If no significant overlap is found, the idea is considered novel and assigned a score of 1.
LLM with guideline: Utilize the NeurIPS 2024 review guidelines to assist LLM in evaluating the novelty of ideas(Table LABEL:tab:LLM_with_guideline_prompt). The final score is determined based on the "Overall" score provided in the review assessment.
LLM with tournament: First, the idea is transformed into the Standardized Project Proposal format(Table LABEL:tab:Standardized_Project_Proposals_prompt). Next, the novelty of all standardized ideas is assessed using the Swiss tournament method, where ideas are iteratively compared in a structured competition. Finally, each idea is assigned a score based on the number of wins it accumulates throughout the tournament.
@@ 4.2 Accuracy Evaluation
As shown in Table 2, our enhanced neighbor density-based novelty measurement algorithm outperforms all baseline models on both the Nature Medicine and Mixed test sets, while also demonstrating strong performance on the NeurIPS test set.
@@
By comparing the results of various LLM-related algorithms, we observe a key similarity between Sonnet-3.7 with guideline and Sonnet-3.7 with tournament: both methods provide very limited external knowledge to the LLM, with no existing literature being fed into the model. As a result, the model’s judgment of novelty is highly inaccurate. In contrast, the LLM + literature search method inputs the 10 most relevant papers to the idea, significantly improving the accuracy of the model’s judgment. Moreover, the accuracy of the LLM + literature search method is much higher in the field of computer science compared to the field of biomedicine, highlighting the significant impact of the model’s internal knowledge on the judgment outcomes, even with the addition of external knowledge. Additionally, Sonnet-3.7(Anthropic (2025)) and Deepseek-r1(Guo et al. (2025)) show much higher AUROC scores than GPT-4o, indicating that when external knowledge is provided, the performance of the inference model greatly surpasses that of the autoregressive model.
However, we observed that the Historical Dissimilarity (HD) metric closely matched the performance of our proposed method on the Nature Medicine and NeurIPS test set. In contrast, on the Mixed test set, there was a significant disparity, with our method achieving an AUROC of 0.795, while HD only reached 0.362. This prompted us to further investigate the underlying reasons for this substantial difference.
The score distributions provided by the Historical Dissimilarity (HD) metric on the NeurIPS and Nature Medicine test set, as shown in Figure 1, are markedly different. This disparity implies that some negative samples from NeurIPS would be evaluated as more novel than some positive samples from Nature Medicine under this evaluation system, highlighting HD’s limited generalization ability across domains. In contrast, the score distributions of our method on both test sets are nearly identical, indicating that our scores are absolute and unaffected by the specific discipline or field. This means that our scores are universally comparable across domains. This result underscores the robust cross-domain evaluation capability of our method, making it applicable for researchers in any field.
Figure 1: Comparison of HD & Our score distributions in different domains. 1: In the right panel, the upper and lower bounds of the score exceeded the actual score range ($[0,100]$) because of linear interpolation. 2: to make the horizontal axis comparable, we scaled the Historical Dissimilarity scores by ×$\times 100$.
@@ 4.3.1 Sensitivity of Hyper-Parameter
Since our method has two key parameters: $P$ and $Q$, we conducted experiments to understand the contribution of each parameter to our algorithm.
@@ 4.3.2 Sensitivity of Design
Relative Neighbor Density(Ours) Absolute Local Density Euclidean distance
@@
4.4 Case Study
We visualize the neighbors of both a novel and a non-novel idea in the embedding vector space to demonstrate the superiority of our algorithm. Figures 3 and 4 show the visualization results of the embedding vectors of an idea and its neighbors on a two-dimensional plane, after dimensionality reduction using t-Distributed Stochastic Neighbor Embedding (t-SNE). While t-SNE excels at preserving the local structure of the data, it does not reliably retain the global structure(Van der Maaten and Hinton (2008)). As a result, the distance between the idea and its $P$ adjacent neighbors is not accurately preserved, but distances between these $P$ neighbors and their $Q$ nearest neighbors is well preserved.
We first use Attention is All You Need (Vaswani et al. (2017)), a highly cited article, as a non-novel idea from the current perspective. Figure 3 clearly illustrates that there is a dense cluster of neighbors around the idea. In contrast, the neighbors around the idea’s $P$ neighbors are relatively few and sparse.
Next, we use Evaluating the World Model Implicit in a Generative Model (Vafa et al. (2025)), an article considered highly novel by the NeurIPS 2024 Program Chairs, as an example of a novel idea, based on their comments (openreview (2025)). In Figure 4, it is evident that the idea’s local neighbor density is much sparser than its $P$ nearest neighbors.
The experimental results demonstrate that the novelty of an idea is reflected in the local structure of the most similar documents to the idea within the embedding vector space, which supports the correctness of our algorithm in principle. Furthermore, a key difference between Figures 3 and 4 is that Figure 3 shows multiple neighboring clusters centered around the idea’s $P$ nearest neighbors, suggesting that the vector density of the two images in the embedding space is notably different. This highlights that the novelty of an idea cannot be determined solely by local density of the idea but must also take into account the vector density surrounding the idea. This is also clearly reflected in the experimental results for the Mixed test set in Table 3.

### v3

@@ Abstract
As large language models (LLMs) demonstrate increasing capability in generating research ideas, automating novelty evaluation has become a critical challenge for AI-driven scientific discovery. This paper presents Relative Neighbor Density (RND), a domain-agnostic algorithm for novelty assessment in research ideas that overcomes the limitations of existing approaches by comparing an idea’s local density with its adjacent neighbors’ densities. We first developed a scalable methodology to create test set without expert labeling, addressing a fundamental challenge in novelty assessment. Using these test sets, we demonstrate that our RND algorithm achieves state-of-the-art (SOTA) performance in computer science (AUROC=0.820) and biomedical research (AUROC=0.765) domains. Most significantly, while SOTA models like Sonnet-3.7 and existing metrics show domain-specific performance degradation, RND maintains consistent accuracies across domains by its domain-invariant property, outperforming all benchmarks by a substantial margin (0.795 v.s. 0.597) on cross-domain evaluation. These results validate RND as a generalizable solution for automated novelty assessment in scientific research.
@@ 1 Introduction
As AI research advances toward AGI, automating scientific discovery processes becomes increasingly important. While large language models (LLMs) now demonstrate capability in generating research ideas, such as AI scientist (Lu et al. (2024)), a critical challenge remains in how to pick pearls from those generated ideas, i.e., how to reliably evaluate these ideas for novelty. Traditionally, novelty in research ideas has been assessed through expert evaluation, where domain specialists judge originality based on their knowledge and experience. However, such assessments are inherently subjective, time-consuming, and inconsistent across evaluators. Automated methods for novelty assessment are therefore crucial for identifying truly innovative directions, particularly as AI systems begin to participate more actively in the research process.
Existing approaches primarily fall into two categories: (1) leveraging large language models (LLMs) as judges and (2) using absolute local density-based novelty metrics. LLM-based approaches include Swiss system tournaments (Si et al. (2024); Hu et al. (2024)) and augmented judgment using NeurIPS review guidelines and Semantic Scholar API (Lu et al. (2024); Su et al. (2024)). Alternatively, absolute local density metrics measure novelty through Euclidean distances between embeddings, as demonstrated by Su et al.’s Overall Novelty (ON) metric (Su et al. (2024)), which combines Historical Dissimilarity, Contemporary Dissimilarity, and citation-based Contemporary Impact, showing strong correlation with human-labeled novelty.
Despite these advances, significant challenges remain. First of all, in studies of any existing approach, the validation methodologies typically rely on small, domain-specific test sets with human-labeled data that quickly become outdated as research advances, limiting scalability and long-term accuracy. Additionally, the LLM-based judgment exhibits sensitivity to input perturbations (Zhuo et al. (2023); Singh et al. (2024)), potentially yielding inconsistent novelty ratings for semantically identical ideas. Finally, absolute local density metrics suffer from arbitrary parameter choices and lack generalizability across research domains with varying citation patterns and publication velocities.
@@
To address these challenges, we establish comprehensive semantic embedding databases for novelty assessment, incorporating publications from two distinct domains: Pubmed, the leading biomedical literature search engine with nearly 36 million articles (Jin et al. (2024)), and Arxiv, which contains more than 2.3 million scholarly articles across eight subject areas (Cornell Tech (2023)). Based on these resources, we develop a methodology to create test sets with trustworthy novelty labels without requiring expert manual labeling.
@@
Furthermore, we propose the Relative Neighbor Density (RND) algorithm, which measures novelty by analyzing the distribution patterns of semantic neighbors rather than simple absolute local density (Figure 1). This approach proves more reliable than LLM-based judgments and more generalizable than existing absolute local density-based metrics across different research domains. Our extensive evaluations using test sets from computer science, biomedical science, and cross-domain contexts demonstrate that our proposed algorithm maintains accuracy within specific domains while scaling effectively across diverse research areas.
@@
Figure 1: Illustration of the Relative Neighbor Density (RND) algorithm. A1/B1: In this step, both the given idea (triangle in A1 and pentagon in B1) and all existing literature in a given research domain are represented in a semantic embedding space. The $P$ nearest neighbors (A/B/C or A’/B’/C’) of the given idea are identified. Then, for each of these neighbors, the neighbor density is computed by identifying $Q$ nearest surrounding neighbors (neighbor sets s1-s3 in A1 and s1’-s3’ in B1). A2/B2: The neighbor densities of the $P$ closest pieces of literature and the given idea are sorted. The RND score of the given idea is determined based on its relative rank among these neighbor densities.
@@ 2.2 Absolute Local Density-based Metrics
An alternative approach focuses on semantic local density to evaluate novelty. Su et al. (Su et al. (2024)) used the Historical Dissimilarity (HD), which is the average Euclidean distance between the generated abstract embedding and embeddings of the 5 most similar abstracts in the historical literature base. We denote it as "Absolute Local Density" because the average distance is a metric of local density, and they use the value of density directly.
In addition to HD, they also developed Overall Novelty (ON) as below, where CI is Contemporary Impact, the average citation count of the top 5 most similar abstracts in contemporary literature base; CD is Contemporary Dissimilarity, calculated by same algorithm in contemporary literature base.
@@ 3.2 Semantic Embedding & Literature Database
For each paper, two semantic embedding vectors were generated—one from its title and another from its abstract—using the M3-Embedding model Chen et al. (2024). The embedding vector dimension, denoted as dimsdims, is 1024. All texts and embedding vectors were stored in Elasticsearch Version 8 for efficient retrieval.
@@ 3.3 Algorithm
In Appendix A, the pseudocode for ND calculation is presented in Algorithm 1; and the pseudocode of complete algorithm is provided in Algorithm 2.
In this algorithm, the selection of values of $P$ and $Q$ is a trade-off between reliability of estimation and other cost. Refer to Appendix A for mathematical analysis and Section 4.3.1 for empirical experimental results. Based on analysis and empirical experiments, we set $P=100$ and $Q=50$.
@@ 4.1.1 Test Set
Mixed Test set: The mixed test set is the combination of the negative samples from NeurIPS and positive samples from Nature Medicine.
@@ 4.1.2 Baseline
To evaluate our algorithm, we selected all existing novelty assessment algorithms as baselines, categorized into two groups: LLM-based and non-LLM-based. Non-LLM-based algorithms, including Relative Neighbor Density(Ours), Historical Dissimilarity(HD), and Overall Novelty(ON), rely solely on literature search and mathematical calculations. Since the output of the literature search for the same query remains consistent, we conducted a single test to assess the algorithm’s performance. In contrast, for LLM-based algorithms, due to the inherent variability of LLM outputs, we ran three tests for each algorithm, calculated the average result, and included the standard deviation in the results.
For all algorithms tested, we use the abstracts of the papers as "ideas" for testing.
Historical Dissimilarity (HD) & Overall Novelty (ON): Both HD and ON are described in Section 2.2. During calculation, the historical database contains papers from 2011 to 2021, and the contemporary database contains papers from 2021 to 2025.
LLM + literature search: Provide LLM with the titles and abstracts of the 10 most relevant papers to the given idea. The model then assesses whether the core concepts of these papers significantly overlap with the idea(Table LABEL:tab:LLM_with_paper_prompt). Sonnet-3.7 (Anthropic (2025)), Deepseek-r1 (Guo et al. (2025)) and GPT-4o (Hurst et al. (2024)) were selected as the LLMs to be tested.
LLM with guideline: Utilize the NeurIPS 2024 review guidelines to assist LLM in evaluating the novelty of ideas(Table LABEL:tab:LLM_with_guideline_prompt). Sonnet-3.7 is selected as the LLM to be tested.
LLM with tournament: First, the idea is transformed into the Standardized Project Proposal format(Table LABEL:tab:Standardized_Project_Proposals_prompt). Next, the novelty of all standardized ideas is assessed using the Swiss tournament method, where ideas are iteratively compared in a structured competition. Finally, each idea is assigned a score based on the number of wins it accumulates throughout the tournament. Sonnet-3.7 is selected as the LLM to be tested.
@@ 4.2 Accuracy Evaluation
As shown in Table 2, our enhanced neighbor density-based novelty measurement algorithm outperforms all baseline models on the Nature Medicine and Mixed test sets, and performs strongly on the NeurIPS test set.
By comparing the results of various LLM-related algorithms, we find that both Sonnet-3.7 with guideline and Sonnet-3.7 with tournament show limited external knowledge input, leading to inaccurate novelty judgments. In contrast, the LLM + literature search method, which incorporates the 10 most relevant papers based on semantic embedding, significantly improves accuracy. This method shows a higher AUROC in computer science than in biomedicine, highlighting the impact of the model’s internal knowledge on judgment outcomes.
Additionally, Sonnet-3.7 and Deepseek-r1 outperform GPT-4o, suggesting that external knowledge enhances the performance of the inference model compared to autoregressive models.
We also observed that the Historical Dissimilarity (HD) metric performed similarly to our method on the Nature Medicine and NeurIPS test sets. However, on the Mixed test set, there was a significant disparity, with our method achieving an AUROC of 0.795, while HD reached only 0.362. This difference reflects the limited generalization ability of HD across domains. In contrast, the distributions of our method’s scores are consistent across all test sets, indicating its robust cross-domain evaluation capability, making it applicable to various research fields.
Figure 2: Comparison of HD & Our score distributions in different domains. 1: In the right panel, the upper and lower bounds of the score exceeded the actual score range ($[0,100]$) because of linear interpolation. 2: to make the horizontal axis comparable, we scaled the Historical Dissimilarity scores by $\times 100$.
@@ 4.3.2 Sensitivity of Design
Ours Absolute Local Density Euclidean distance
@@
4.4 Case Study
To visualize the geometric structure of ideas and their neighboring points in the embedding space, we selected two ideas and plotted their surrounding neighbors on a 2-D plane using t-SNE (Van der Maaten and Hinton (2008)).
We first examine Attention is All You Need (Vaswani et al. (2017)), a highly cited article, as an example of a non-novel idea from the current perspective. As shown in Figure 4, there is a dense cluster of neighbors surrounding the idea, while the neighbors around its $P$ nearest neighbors are relatively fewer and more sparsely distributed.
Next, we analyze Evaluating the World Model Implicit in a Generative Model (Vafa et al. (2025)), an article recognized as highly novel by the NeurIPS 2024 Program Chairs based on their comments (openreview (2025)). In Figure 5, the idea’s local neighborhood appears considerably sparser than that of its $P$ nearest neighbors.
Moreover, the differences between the two figures extend beyond the local density around the central idea; the local densities of neighboring ideas also exhibit distinct patterns. In Figure 4, multiple neighboring clusters centered around the idea’s $P$ nearest neighbors are visible, whereas such clustering patterns are absent in the novel idea’s neighborhood. These observations suggest that an idea’s novelty is associated not only with its own local density but also with the density distribution of its neighboring ideas.